OpenStudy (anonymous):

Find the numerical value of cosh (ln5)

hartnn (hartnn):

know the definition of cosh x ?

hartnn (hartnn):

no?

OpenStudy (anonymous):

yyes. e^x - e^-2 /2

OpenStudy (anonymous):

e^-x *

hartnn (hartnn):

now use ,the fact that \(\huge e^{ln a}=a\)

OpenStudy (anonymous):

sorry it is + sign in between the e

hartnn (hartnn):

yesm i know, now just put x=ln 5

OpenStudy (anonymous):

i did, but i dont know what ot do for e^ -ln 5

OpenStudy (anonymous):

where did that -ve sign go..

OpenStudy (anonymous):

cuz it will become 5 + (-5) / 2 , then asnwer will be 0

hartnn (hartnn):

for that , use log property \(\large n log a=log a^n\) n=-1 here

hartnn (hartnn):

so -ln 5 will be ?

OpenStudy (anonymous):

so it is loge^-ln4?

hartnn (hartnn):

4 ?? how u got 4 ?

OpenStudy (anonymous):

sorry 5, typo

OpenStudy (anonymous):

original is e^-ln5 -ve is not the power of ln5, so how can u use log to bring it down?

hartnn (hartnn):

\(\large n log a=log a^n \\ \large -log 5 =log (5^{-1})\) got this ?

OpenStudy (anonymous):

log = ln?

hartnn (hartnn):

now use \(\large e^{ln (5^{-1})}=?\) yes, its ln

OpenStudy (anonymous):

5 + 1/5 ) /2 26/5 / 2= 13/5

hartnn (hartnn):

thats correct, did u understand ?

OpenStudy (anonymous):

yes i do, thanks!

hartnn (hartnn):

welcome ^_^